# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ V Toward Efficient Photonic AI Computing: A Software-hardware Co-design Perspective
In this review, we have explored the promising potential of photonics for AI acceleration. As a highly interdisciplinary field, photonic AI computing requires contributions from various domains, including device, circuit, architecture, and algorithm levels. Previous sections have examined a range of efforts from academia and industry from device, circuit, and architecture parts. However, solely focusing on progress at the individual component level is insufficient to fully unlock the vast potential of photonic computing. Instead, a holistic approach across multiple levels is crucial for maximizing the performance of photonic AI systems, necessitating a software-hardware co-design perspective.
@@ V-A2 Co-design progress
The concern over area cost has prompted the development of compact photonic devices and the advancement of fabrication processes. This, combined with various hardware-efficient PTC designs discussed in Section III-B3, aims to address the challenge of large spatial footprints in photonic integrated circuits. Beyond the methods focusing solely on device or circuit levels, researchers are exploring holistic device-circuit-algorithm co-design efforts by exploring two promising directions: (1) domain-specific photonic computing engines to trade off between matrix expressivity and hardware efficiency; (2) AI-assisted automatic compact PTC design.
@@
The first line of work focuses on building developing domain-specific photonic computing engines instead of universal linear units in previous works [157, 158, 81, 115, 159, 117, 148], such as MZI mesh and MRR bank [29, 110]. The over-parameterization of neural networks has inspired various research efforts exploring the construction of efficient neural networks beyond conventional GEMM with restricted matrix parameter space, including low-rank neural networks and structured neural networks [160, 161]. We refer to such neural networks with restricted weight matrix space as subspace neural networks [81]. Subspace neural networks have shown considerable improvements in efficiency while maintaining comparable representability to traditional neural networks. The success of efficient subspace neural networks can be leveraged in ONNs by sacrificing the universality of weight representation in exchange for higher hardware efficiency. For instance, Gu et al. proposes to implement the efficient circulant neural network and devise a novel butterfly photonic architecture with improved area efficiency over previous circuits[157, 161]. It implements the optical fast Fourier transform (OFFT) and its inverse (OIFFT), which are used to efficiently perform circulant matrix multiplication. Gu et al. further extend the proposed architecture to a trainable transform structure to enable the implementation of more matrix transformation [158]. Moreover, to solve the issue of quadratic increase in the number of MZI devices when supporting larger matrices in MZI mesh [29], Xiao et al. apply tensor-train decomposition first to decompose large over-parameterized weight matrices into smaller ones, thus substantially reducing the number of MZI devices required [159].
@@
Another noteworthy advancement in this field is the exploration of automatic PTC design, departing from the conventional manual design paradigm. In this approach, the footprint can be incorporated as a constraint to enable the automatic generation of compact PTC. For instance, ADPET introduces the first automatic AI-assisted differentiable search framework for PTC topology design [162]. It first constructs a probabilistic photonic SuperMesh and then employs differentiable optimization in a huge and highly discrete PTC search space. This framework adapts to various circuit footprint constraints and foundry PDKs. The PICs developed using this method demonstrate a substantial increase in footprint compactness, ranging from 2 to 30$\times$ compared to both the MZI mesh and the manually designed compact butterfly mesh. This automated approach promises to revolutionize the design process, enabling the creation of more efficient and compact PTCs for AI applications.
@@ V-B1 Issue
Energy efficiency is a crucial metric when evaluating computing hardware. To evaluate the energy efficiency of photonic AI hardware, given their mixed-signal setup, it’s crucial to consider the energy costs associated with both digital and optical parts of the system. Following previous studies [114, 163], the energy cost of photonic computing can be broadly categorized into optical costs associated with performing computation and electrical costs related to loading operands ($X$ and $Y$) and detecting output ($O$), which can be expressed as follows:
@@
Here, $E_{\text{load}}$ encompasses the energy costs associated with memory reading ($E_{\text{read}}$), D/A conversion ($E_{\text{DAC}}$), and signal modulation ($E_{\text{mod}}$). $E_{\text{det}}$ represents the costs of optical signals detection ($E_{\text{PD}}$), signal amplification ($E_{\text{amp}}$), A/D conversion ($E_{\text{ADC}}$), and the subsequent writing of results back to memory ($E_{\text{write}}$). $E_{\text{comp}}$ includes other energy costs associated with performing computation, which could be negligible for a fully passive PIC.
Among all the components, the transition between digital and analog signals presents a significant bottleneck in the system energy consumption, particularly evident in data movement (memory-associated cost) and ADC/DAC, as labeled in bold in Eq. (19). This transition can occupy more than 80% of the overall energy consumption [150, 164].
@@ V-B2 Co-design Progress
Device-level advancement promises straightforward reduction in energy costs in Eq. (19), such as the progress of energy-efficient ADC [165], efficient optical modulators [166, 167], and on-chip laser [168]. However, in this section, we explore recent advancements beyond the device level aimed at improving energy efficiency, with a particular focus on addressing signal conversion and memory aspects.
At the circuit and architecture levels, several efforts have been made to reduce signal conversion and data movement costs, as summarized in Table II. Optical broadcast [110] is a widely-used technique that enables spatial sharing of encoded signals, thereby saving on DAC and E-O modulation costs. This approach has been further extended in a crossbar-style design in Ref [114] to enable both input and weight to be spatially shared. Another strategy involves keeping weights static in photonic devices or employing non-volatile devices to reuse encoded signals across different inputs temporally. Additionally, leveraging spectral parallelism allows for sharing operands among different inputs. Both approaches can save the encoding costs related to DAC and modulation. Time-integration techniques have been employed to explore analog-domain temporal accumulation [148, 114, 117], significantly reducing the A/D conversion frequency and preserving more computations within the analog domain. Recently, the use of electro-optic analog memory has gained attention [151], where analog memory is placed near photonic devices, and DACs are reused across rows of analog memory devices, thereby reducing DAC costs.
[…]
Method Input X Weight Y Ouput O
[…]
At the algorithm level, many studies employ low-bit quantization methods to maintain accuracy with low-precision weights, inputs, and activations [135, 157, 170, 116]. This can reduce the required DAC and ADC costs with lower resolution, as well as reducing the memory data movement cost. Additionally, the residue number system has been explored to reduce precision requirements by achieving high-precision computing with low-precision ones [171].
@@ V-C1 Issue
Ensuring the correctness of functionality is a basic requirement of computing hardware. However, operating as an analog computing platform, photonic computing hardware inherently grapples with robustness issues stemming from two primary factors, as depicted in Fig. 12: (1) various non-ideal conditions. These conditions include process variations, device noises, environmental factors, and limited endurance. (2) limited precision of inputs and outputs. This limitation arises due to the finite control resolution and the substantial overhead associated with high-precision DACs and ADCs. Unlike low-precision electronics digital computing, as shown in Fig. 12.a, photonic computing uniquely suffers from output precision loss due to the usage of ADC to convert analog outputs back to the digital domain, where the ADC precision is typically unsatisfactory compared to the output precision. To understand the mismatch between ADC and output precision, consider a straightforward scenario in which a dot product computation of length-$N$ is executed by a photonic computing engine. Note that here we consider an ideal case where the precision constraints are on the weights and activations, and the photonic computing engine can be treated as a dot-product engine based on a scalar product. [^1] The real case can be more intricate. For example, low precision may be reflected in device control variables, and photonic computing cannot be easily treated based on scalar product. The multiplication of a $B_{w}$-bit signed weight by a $B_{x}$-bit signed activation results in a product with $B_{w}+B_{x}-2+\log_{2}N$ magnitude bits and one sign bit. Here, the additional $log_{2}N$ bits of precision stem from the addition of $N$ scalar products. However, achieving this desired precision often exceeds the capabilities of affordable ADCs, typically within 8 bits [172]. For instance, a 4-bit weight and 4-bit activation in a length-$32$ dot product computation would ideally yield an output precision of $13$ bits, 5 bits more than 8 bits. This distinction underscores a notable difference between photonic computing and traditional low-precision arithmetic units like int8 GPU tensor cores where outputs are preserved in high precision [173], raising a unique precision issue.
@@ V-C2 Co-design Progress
Here, we discuss the efforts to mitigate precision and the impact of noise on accuracy. For precision issues, previous efforts have often drawn inspiration from quantization, either employing Quanztiation-aware Training (QAT) [135, 174, 170, 116, 175], or Post-training Quantization (PTQ) to tackle input and weight precision issues [176]. These approaches enable neural networks to inherently tolerate low-precision arithmetic, thereby mitigating the impact of precision limitations. Specifically, Gu et al. first developed the QAT flow by directly optimizing low-precision device control signals in the discretized space [135]. However, the aforementioned quantization techniques are insufficient to compensate for the computation information loss due to the low-precision ADC. This is because errors occur in each partial result, and they finally accumulate when we tile a large matrix workload. Some work in analog computing specifically tuning the ADC reference voltage to balance the dynamic quantization range-resolution tradeoff [177]. Furthermore, weight and input slicing are proposed as strategies to manage the trade-off in ADC resolution [178].
@@
For noise mitigation, noise-aware training has emerged as a widely used technique to enhance resilience [134, 135, 120]. This approach involves introducing noises and variations during the training process to enhance the noise resilience of ONNs. Previous studies have focused on modeling various noise sources [179, 120], such as dynamic noise, static manufacturing variation, and thermal crosstalk, and incorporating them into the training process. Furthermore, explicit robustness optimization terms can be integrated into the training process to further enhance robustness against noise. For instance, Ref [135] estimates the noise sensitivity of weights and applies protective regularization terms to sensitive weights during optimization. Similarly, Zhu et al. introduce an additional regularization term on the phase magnitude in MZI mesh to reduce thermal and crosstalk concerns [180]. Additionally, knowledge distillation strategies, as employed in Ref [181], can be employed to guide the optimization of noisy student ONN models under the guidance of a noise-free teacher model. This approach significantly enhances the robustness of models against static process variation and dynamic input signal noises.
@@
However, noise patterns can be highly intricate, making it challenging and time-consuming to analyze and model them accurately. In response to this challenge, some studies have explored training ONNs directly with noisy physical responses from real chips, eliminating the need for explicit noise modeling [81, 136, 137]. This approach, known as physical neural network training, is introduced in detail in Section III-D. Moreover, on-chip training is another way to recover accuracy by inherently modeling on-chip noise during training, which is further explored in Section V-D.
@@
Besides exploring the inherent noise tolerance of neural networks, another straightforward direction is to compress the on-chip noise levels. Specifically, some work [152, 182] advocate employing device-level and circuit-level design space exploration to mitigate the impact of process imperfections and crosstalk effects. ELight [174] suggests avoiding frequent reprogramming on PCM cells by enhancing the similarity of mapping weights to mitigate early wear-out. Additionally, reducing the number of active devices on the circuits can also reduce on-chip noise, as noise-induced errors are generally positively related to the number of noise sources. Pruning redundant devices [162], or weight blocks [158, 157, 81], brings significant noise robustness improvement.
@@ V-D1 Issue
The adaptability or trainability of photonic analog computing platforms is another major challenge in their practical application. The significance of self-learnability arises from three main aspects, as shown in Fig. 13. Firstly, on-chip training empowers optical hardware with high adaptability when the working environment or conditions drift or when workloads switch. Secondly, the realization of on-chip training can unlock many important edge learning applications such as local online learning, transfer learning, and lifelong learning. Moreover, as discussed in Section III-D, on-chip training can address robustness issues in-situand bridge the gap between simulation and real deployment.
@@ V-D2 Co-design Progress
We summarize recent on-chip training protocols, either validated through simulation or real demonstration, in Table III. Generic gradient-free optimization methods, e.g., evolutionary algorithms and brute-force device tuning [183, 29, 105], have been used to optimize on-chip parameters with no gradient information involved. Nevertheless, these approaches encounter difficulties when scaling to large-scale models with slow convergence and poor stability. Besides gradient-free approach, the investigation into the direct feedback alignment (DFA) training algorithm is further explored in [184] for in-situ training [189, 190]. DFA propagates errors through fixed random feedback projections from the output layer to each hidden layer in parallel, thereby obviating the necessity for the sequential backpropagation of gradients. Another approach, the adjoint variable method, was introduced to perform on-chip backpropagation by computing the gradients in-situ with per-device optical field monitors [138], and recent work has experimentally demonstrated the concept [139]. However, scalability remains a significant concern due to considerable hardware support overhead, such as the need for per-device monitoring.
[…]
Figure 13: Self-learnable ONN paradigm exemplified by a one scalable ONN on-chip training framework with in-situ gradient calculation [188].

### v2

@@ V Toward Efficient Photonic AI Computing: A Software-hardware Co-design Perspective
In this review, we have explored the promising potential of photonics for AI acceleration. As a highly interdisciplinary field, photonic AI computing requires contributions from various domains, including device, circuit, architecture, and algorithm levels. Previous sections have examined a range of efforts from academia and industry across the device, circuit, and architecture parts. However, solely focusing on progress at the individual component level is insufficient to fully unlock the vast potential of photonic computing. Instead, a holistic approach across multiple levels is crucial for maximizing the performance of photonic AI systems, necessitating a software-hardware co-design perspective.
@@ V-A2 Co-design progress
The concern over area cost has prompted the development of compact photonic devices and the advancement of fabrication processes. This, combined with various hardware-efficient PTC designs discussed in Section III-B3, aims to address the challenge of large spatial footprints in PICs. Beyond the methods focusing solely on device or circuit levels, researchers are exploring holistic device-circuit-algorithm co-design efforts in two promising directions: (1) domain-specific photonic computing engines to trade off between matrix expressivity and hardware efficiency; (2) AI-assisted automatic compact PTC design.
@@
The first line of focuses on developing domain-specific photonic computing engines [165, 89, 127, 166, 129, 158], instead of universal linear units as in previous works, such as MZI meshes and MRR banks [28, 121]. The over-parameterization of neural networks has inspired various research efforts exploring the construction of efficient neural networks beyond conventional GEMM within a restricted matrix parameter space, including low-rank neural networks and structured neural networks [167, 168]. We refer to such neural networks with a restricted weight matrix space as subspace neural networks [89, 125]. Subspace neural networks have shown considerable improvements in efficiency while maintaining comparable representability to traditional neural networks. The success of efficient subspace neural networks can be leveraged in ONNs by sacrificing the universality of weight representation in exchange for higher hardware efficiency. For instance, Gu et al. proposed to implement an efficient circulant neural network and devise a novel butterfly photonic architecture with improved area efficiency over previous circuits[165, 168]. It implements the optical fast Fourier transform (OFFT) and its inverse (OIFFT), which are used to efficiently perform circulant matrix multiplication. Gu et al. further extended the proposed architecture to a trainable transform structure to enable the implementation of more matrix transformation [165]. Moreover, to solve the issue of quadratic increase in the number of MZI devices when supporting larger matrices in an MZI mesh [28], Xiao et al. applied tensor-train decomposition first to decompose large over-parameterized weight matrices into smaller ones, thus substantially reducing the number of MZI devices required [166].
@@
Another noteworthy advancement in this field is the exploration of automatic PTC design, departing from the manual design paradigm. In this approach, the footprint can be incorporated as a constraint to enable the automatic generation of compact PTC. For instance, ADPET introduces the first automatic AI-assisted differentiable search framework for PTC topology design [169]. It first constructs a probabilistic photonic SuperMesh and then employs differentiable optimization in a huge and highly discrete PTC search space. This framework adapts to various circuit footprint constraints and foundry PDKs. The PICs developed using this method demonstrate a substantial increase in footprint compactness, ranging from 2 to 30$\times$ compared to both the MZI mesh and the manually designed compact butterfly mesh. This automated approach promises to revolutionize the design process, enabling the creation of more efficient and compact PTCs for AI applications.
@@ V-B1 Issue
Energy efficiency is a crucial metric for evaluating computing hardware. When assessing the energy efficiency of photonic AI hardware, especially given its mixed-signal setup, it is essential to consider the energy costs associated with both the digital and optical components of the system. Following previous studies [126, 170], the energy cost of photonic computing can be broadly categorized into optical costs associated with performing computation and electrical costs related to loading operands ($X$ and $Y$) and detecting output ($O$), which can be expressed as follows:
@@
Here, $E_{\text{load}}$ encompasses the energy costs for memory reading ($E_{\text{read}}$), D/A conversion ($E_{\text{DAC}}$), and signal modulation ($E_{\text{mod}}$). $E_{\text{det}}$ represents the costs of optical signal detection ($E_{\text{PD}}$), signal amplification ($E_{\text{amp}}$), A/D conversion ($E_{\text{ADC}}$), and the subsequent writing of results back to memory ($E_{\text{write}}$). $E_{\text{comp}}$ includes other energy costs associated with performing computation, which could be negligible for a fully passive PIC.
Among all components, the transition between digital and analog signals presents a significant bottleneck in the system energy consumption, particularly evident in data movement (memory-associated cost) and ADC/DAC, as labeled in bold in Eq. (21). This transition can occupy more than 80% of the overall energy consumption [160, 171], which is different from the power composition of the optical digital computing outlined in Section II-B.
@@ V-B2 Co-design Progress
Device-level advancement promises straightforward reduction in energy costs in Eq. (21), such as the progress in energy-efficient ADCs [172], efficient optical modulators, and on-chip laser [173]. However, in this section, we explore recent advancements beyond the device level, with a particular focus on addressing the signal conversion and memory aspects.
At the circuit and architecture levels, several efforts have been made to reduce signal conversion and data movement costs, as summarized in Table II. Optical broadcast is a widely-used technique that enables spatial sharing of encoded signals [121], thereby saving on DAC and E-O modulation costs. This approach has been further extended to a crossbar-style design in Ref [126] to enable both input and weight to be spatially shared. Another strategy involves keeping weights static in photonic devices or employing non-volatile devices to reuse encoded signals across different inputs temporally. Additionally, leveraging spectral parallelism allows for sharing operands among different inputs. Both approaches can lower the encoding costs related to DAC and modulation. Time-integration techniques have been employed to explore analog-domain temporal accumulation [132], significantly reducing the A/D conversion frequency and preserving more computations within the analog domain. Recently, the use of E-O analog memory has gained attention [161], where analog memory is placed near photonic devices, and DACs are reused across rows of analog memory, thereby reducing DAC costs.
TABLE II: Comparison of recent approaches on saving signal conversion cost, including D/A ($\textit{E}_{\text{DAC}}$), A/D($\textit{E}_{\text{ADC}}$), E-O($\textit{E}_{\text{mod}}$) energy costs.
Method Input X Weight Y Output O
Optical broadcast [121] $\textit{E}_{\textbf{DAC}},\textit{E}_{\textbf{mod}}\downarrow$
Weight-static dataflow [159] - $\textit{E}_{\textbf{DAC}},\textit{E}_{\textbf{mod}}\downarrow$
Spectral parallelism [124, 157] - $\textit{E}_{\textbf{DAC}},\textit{E}_{\textbf{mod}}\downarrow$
Time-integration [158, 126, 129] - - $\textit{E}_{\textbf{ADC}}\downarrow$
Electro-Optic analog memory [161] - $\textit{E}_{\textbf{DAC}}\downarrow$
Non-volatile device [124, 174] - $\textit{E}_{\textbf{mod}}\downarrow$
At the algorithm level, many studies have adopted low-bit quantization techniques to preserve accuracy while using low-precision weights, inputs, and activations. [146, 165, 175, 128]. This approach can reduce the required DAC and ADC costs with lower resolution, as well as reduce the memory data movement cost. Furthermore, the residue number system has been explored as a method to reduce precision requirements by achieving high-precision computations using low-precision components [176].
@@ V-C1 Issue
Ensuring functional correctness is a fundamental requirement of computing hardware. However, analog photonic computing inherently faces robustness challenges due to two primary factors, as depicted in Fig. 12: (1) Various non-ideal conditions include process variations, device noises, environmental factors, and limited endurance. (2) Limited precision of inputs and outputs arises from the finite control resolution and the substantial overhead associated with high-precision DACs and ADCs. Unlike low-precision electronics digital computing, as shown in Fig. 12.a, photonic computing uniquely suffers from output precision loss in the conversion of analog outputs back to the digital domain, where the ADC precision is typically unsatisfactory compared to the output precision. To understand the mismatch between an ADC and output precision, consider a straightforward scenario in which a photonic computing engine executes a dot product operation of length $N$. Note that here we consider an ideal case where the precision constraints are on the weights and activations, and the photonic computing engine can be treated as a dot-product engine based on a scalar product. [^1] The real case can be more intricate. For example, low precision may be reflected in device control variables, and photonic computing cannot be easily treated based on scalar product. The multiplication of a $B_{w}$-bit signed weight by a $B_{x}$-bit signed activation results in a product with $B_{w}+B_{x}-2+\log_{2}N$ magnitude bits and one sign bit. Here, the additional $log_{2}N$ bits of precision stem from the addition of $N$ scalar products. However, achieving this desired precision often exceeds the capabilities of affordable ADCs, typically within 8 bits [177]. For instance, a 4-bit weight and 4-bit activation in a length-$32$ dot product computation would ideally yield an output precision of 13 bits. The 5-bit overhead underscores a notable difference between photonic computing and traditional low-precision arithmetic units like int8 GPU tensor cores where outputs are preserved in high precision [178], raising a unique precision issue.
@@ V-C2 Co-design Progress
Here, we explore efforts to mitigate the precision issue and the impact of noise on accuracy. Regarding precision issues, previous efforts have often drawn inspiration from quantization, either employing Quanztiation-aware Training (QAT) [146, 179, 175, 128, 180], or Post-training Quantization (PTQ) [181]. These approaches enable neural networks to inherently tolerate low-precision arithmetic, thereby mitigating the impact of precision limitations. Specifically, Gu et al. first developed the QAT flow by directly optimizing low-precision device control signals in the discretized space [146]. However, the aforementioned quantization techniques are insufficient to compensate for the computation information loss due to the low-precision ADC. This is because errors occur in each partial result, and they finally accumulate when we tile a large matrix workload. Some work in analog computing specifically tune the ADC reference voltage to balance the tradeoff between dynamic quantization range and solution [182]. Furthermore, weight and input slicing are proposed as strategies to manage the trade-off in ADC resolution [183].
@@
For noise mitigation, noise-aware training has emerged as a widely used technique to enhance resilience [145, 146, 132]. This approach involves introducing noises and variations during the training process to enhance the noise resilience of ONNs. Previous studies have focused on modeling various noise sources [184, 132], such as dynamic noise, static manufacturing variation, and thermal crosstalk, then incorporated them into the training process. Furthermore, explicit robustness optimization terms can be integrated into the training process to further enhance robustness against noise. For instance, Ref [146] estimates the noise sensitivity of weights and applies protective regularization terms to sensitive weights during optimization. Similarly, Zhu et al. introduced an additional regularization term on the phase magnitude in an MZI mesh to reduce the crosstalk concerns [185]. Additionally, knowledge distillation strategies, as employed in Ref [186], can be employed to guide the optimization of noisy student ONN models under the guidance of a noise-free teacher model. This approach significantly enhances the robustness of models against static process variation and dynamic input signal noises.
@@
However, noise patterns can be highly intricate, rendering the analysis and modeling of these patterns both challenging and time-consuming. In response to this challenge, some studies have explored training ONNs directly with noisy physical responses from real chips, eliminating the need for explicit noise modeling [89, 147, 148]. This approach, known as physical neural network training, has been introduced in detail in Section III-D. Moreover, on-chip training is another way to recover accuracy by inherently modeling on-chip noise during training, which is further explored in Section V-D.
@@
Besides exploring the inherent noise tolerance of neural networks, another straightforward direction is to compress the on-chip noise levels. Specifically, some work advocate employing device-level and circuit-level design space exploration to mitigate the impact of process variation and crosstalk [187, 188]. Ref [179] suggests avoiding frequent reprogramming on PCM cells by enhancing the similarity of mapping weights to mitigate early wear-out. Additionally, reducing the number of active devices in PICs can also decrease on-chip noise, as noise-induced errors typically correlate positively with the number of noise sources. Therefore, pruning redundant devices [169], or weight blocks [165, 89], brings significant noise robustness improvement.
@@ V-D1 Issue
The adaptability or trainability of photonic analog computing platforms is another major challenge in their practical application. The significance of self-learnability arises from three main aspects, as shown in Fig. 13. Firstly, on-chip training enhances the adaptability of optical hardware when working conditions drift, or when workloads change. Secondly, the realization of on-chip training can unlock many important edge learning applications such as local online learning, transfer learning, and lifelong learning. Moreover, as discussed in Section III-D, on-chip training can address robustness issues in-situ and bridge the gap between simulation and real implementation.
@@ V-D2 Co-design Progress
We summarize recent on-chip training protocols, either validated through simulation or real demonstration, in Table III. Generic gradient-free optimization methods, e.g., evolutionary algorithms and brute-force device tuning [189, 28, 116], have been used to optimize on-chip parameters with no gradient information involved. Nevertheless, these approaches encounter difficulties when scaling to large-scale models with slow convergence and poor stability. Besides the gradient-free approach, investigation into the direct feedback alignment (DFA) training algorithm is further explored in Ref [190] for in-situ training [194, 195]. DFA propagates errors through fixed random feedback projections from the output layer to each hidden layer in parallel, thereby obviating the necessity for the sequential backpropagation of gradients. Another approach, the adjoint variable method, was introduced to perform on-chip backpropagation by computing the gradients in-situ with per-device optical field monitors [150], and recent work has experimentally demonstrated the concept [151]. However, scalability remains a significant concern due to considerable hardware support overhead, such as the need for per-device monitoring.
On-chip training protocols based on forward-only zeroth-order optimization aim to approximate gradients using the finite difference method, perturbing parameters with small random values for gradient estimation. Instead of adjusting individual on-chip parameters separately, Bandyopadhyay et al. proposed a method that samples perturbations for all parameters and shares them among all training instances in a single iteration [191]. Similarly, FLOPS, proposed in Ref [192], shares the sampled perturbations among single mini-batches instead of all mini-batches in the same interaction, showing better convergence than Ref [191]. To further enhance the scalability of zeroth-order training protocols, MixedTrain proposed in Ref [193] partitions PICs into passive and active regions and trains only a small subset of active devices in each iteration, ensuring the capacity to handle a larger scale ONN compared to FLOPS. The zeroth-order-based approach shows the training scale to thousands of MZIs as shown in Ref [192, 193], while it is still not enough when dealing with modern-size machine learning models. L2ight introduces a subspace optimization algorithm and develops a method for in-situ calculation of first-order subspace gradients [149], as shown in Fig. 13. Additionally, a multi-level hardware-aware sparse training method is employed to boost training efficiency. This study demonstrates the first instance of training a million-parameter-level ONN, showcasing exceptional stability.
Figure 13: The self-learnable ONN paradigm exemplified by a single scalable ONN on-chip training framework with in-situ gradient calculation [149].